# Asymptotic utility of spectral anonymization

### v1

@@ 1 Introduction
Figure 1: Illustration of the effects of $\mathcal{P}$-SA (blue), $\mathcal{J}$-SA (orange) and $\mathcal{O}$-SA (purple) to the original data (green), with numerical labels corresponding to the row indices of original data, in a scenario with $n=10$ and $p=2$. $\mathcal{P}$-SA and $\mathcal{J}$-SA occasionally result in unwanted overlap, wherein a row in the anonymized data matches another row in the original dataset. In the case of $\mathcal{J}$-SA, this occurrence is due to coincidental alignment of signs with those in the original singular vectors, causing the observation to be duplicated (overlap of same indices), whereas for $\mathcal{P}$-SA, random permutation of values may cause another row to align with one in the original dataset (overlap of different indices). Conversely, $\mathcal{O}$-SA induces arbitrary rotations in the spectral space that prevent exact matches altogether with probability 1.
As our second main contribution, we investigate the earlier observation, stated also by Lasko and Vinterbo [10], that $\mathcal{P}$-SA approximately preserves the means, variances, covariances, and linear correlations of the original data $X_{n}$. In this work, we quantify this statement in an asymptotic framework, working, for simplicity, under the assumption that the rows of $X_{n}$ are sampled i.i.d. from the normal distribution $\mathcal{N}_{p}(\mu,\Sigma)$ for some fixed $\mu\in\mathbb{R}^{p}$ and positive definite covariance matrix $\Sigma$. Extensions to other distributional assumptions are investigated through simulations in Section 3 and discussed further in Section 4. As the multivariate normal distribution is entirely determined by its means and covariances, evaluating the resemblance of anonymized data with respect to these moments is equivalent to assessing general data utility in the present scenario.
@@ 2 Asymptotic utility
Theorem 2.2 offers us two insights: (a) All three forms of SA are equally efficient in covariance estimation. (b) Assume next that $O=I_{p}$ in the eigendecomposition $\Sigma=O\Lambda O^{\prime}$ (since the rotation $O$ only fixes the coordinate system in which the data are observed, this choice does not limit our interpretations). Then, for instance when $p=2$, the limiting covariance matrices of $S_{n}$ and $S_{n,\mathcal{A}}$ take the forms:
@@ 3 Simulations
Figure 2: The relative error of the empirical covariance matrices of sample mean (solid line) and sample covariance (dotted line) compared to their asymptotic covariance matrices for $\mathcal{P}$-SA (blue), $\mathcal{J}$-SA (orange), and $\mathcal{O}$-SA (purple), when original data (green) is sampled from a normal distribution meeting the Assumption 1. Sample size is presented on a logarithmic scale.
@@ 3.1 Finite-sample utility
Figure 3: The relative error of the empirical covariance matrices of sample mean (solid line) and sample covariance (dotted line) compared to their asymptotic covariance matrices for $\mathcal{P}$-SA (blue), $\mathcal{J}$-SA (orange), and $\mathcal{O}$-SA (purple), when original data (green) is sampled from a normal distribution violating the Assumption 1. Sample size is presented on a logarithmic scale.
@@ 4 Conclusion
Based on our findings with finite data, none of the three methods demonstrated clear superiority in utility convergence. However, concerning privacy, $\mathcal{O}$-SA outperformed $\mathcal{P}$-SA and $\mathcal{J}$-SA by never generating identical records. Hence, we advocate for employing $\mathcal{O}$-SA for data anonymization, notwithstanding its computational intensity compared to $\mathcal{P}$-SA and $\mathcal{J}$-SA. For a dataset size of $n=1000$ and $p=6$, the runtime for $\mathcal{O}$-SA is on average 2.44 seconds, while for $\mathcal{P}$-SA and $\mathcal{J}$-SA, it is 0.002 and 0.001 seconds, respectively (using AMD Ryzen 5 5600X). In practical scenarios, data anonymization is typically a one-time operation, mitigating concerns over computational overhead. However, should computational efficiency become paramount, $\mathcal{P}$-SA emerges as the secondary choice, demonstrating no matches in our simulations with $p=6$. In case privacy is given secondary priority, then $\mathcal{P}$-SA offers superior mean estimation efficiency to its two competitors (Theorem 2.1) and should be used. Nonetheless, further research is warranted to evaluate these SA algorithms using empirical real-world data and utility scenarios.
We conclude with comments regarding our assumptions. Our simulations revealed that Assumption 1 cannot be dispensed with and, indeed, it is relied upon extensively in our proof techniques in Appendix 0.A. And while the case of $\Sigma$ having eigenvalues with multiplicity greater than one has no practical importance, studying also this case would nevertheless complete the theoretical story behind the method. Application to Poisson-distributed data revealed that our results do not directly extend outside of normality, which is to be expected as, again, our proofs use this fact in several places. However, it might be that the results continue to hold if the normal distribution is replaced with another distribution that has the same moment sequence up to a certain moment. This behavior is known as universality in high-dimensional statistics [7] and we leave its study in the current context for future work.
@@ 4.0.1 Acknowledgements
The study was supported by the Finnish Cultural Foundation (grant 00220801) and the Research Council of Finland (grants 347501 and 353769).
@@ Appendix 0.A Proofs of technical results
For $\mathcal{O}$-SA, the proof is similar to $\mathcal{J}$-SA, and we point out only the differences next. To see that $X_{n}^{\prime}H_{n}O_{nk}1_{n}/\sqrt{n}$ is stochastically bounded, we observe that $O_{nk}1_{n}\sim\sqrt{n}O_{nk}e_{1}\sim\sqrt{n}u_{n}$, where $u_{n}$ is uniform on the unit sphere $\mathbb{S}^{n-1}$. Moreover, letting $\chi_{n}$ denote a random variable obeying a $\chi$-distribution with $n$ degrees of freedom that is independent of $u_{n}$, we have $b_{n}:=\chi_{n}u_{n}\sim\mathcal{N}_{n}(0,I_{n})$ and $\chi_{n}/\sqrt{n}\rightarrow_{p}1$. Consequently, denoting a generic column of $X_{n}$ by $x_{n}\sim\mathcal{N}_{n}(0,\lambda I_{n})$, we have $(1/\sqrt{n})x_{n}^{\prime}H_{n}O_{nk}1_{n}\sim x_{n}^{\prime}H_{n}u_{n}=x_{n}^{\prime}u_{n}-\bar{x}_{n}1_{n}^{\prime}u_{n}=b_{n}^{-1}(1/\sqrt{n})x_{n}^{\prime}z_{n}-\sqrt{n}\bar{x}_{n}\bar{z}_{n}b_{n}^{-1}$, showing that $X_{n}^{\prime}H_{n}O_{nk}1_{n}/\sqrt{n}$ is stochastically bounded. Hence,
$(\sqrt{n}\bar{y}_{n})_{k}=$ $(\sqrt{n}\bar{x}_{n})_{k}+b_{n}^{-1}\frac{1}{\sqrt{n}}e_{k}^{\prime}X_{n}^{\prime}z_{n}+o_{p}(1)=(\sqrt{n}\bar{x}_{n})_{k}+\frac{1}{\sqrt{n}}e_{k}^{\prime}X_{n}^{\prime}z_{n}+o_{p}(1)$
$=$ $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}x_{n,ik}(1+z_{n,i})+o_{p}(1).$
@@
since $\sqrt{n}\bar{y}_{n}=\mathcal{O}_{p}(1)$. From the proof of Theorem 2.1 we have that $V_{n}D_{n}U_{n0}^{\prime}1_{n}\bar{x}_{n}/\sqrt{n}=o_{p}(1)$, meaning that
@@
where $G_{n}\in\mathbb{R}^{p\times p}$ is a matrix with the elements $g_{n,kk}=0$, $g_{n,k\ell}=u_{nk}^{\prime}T_{nk}^{\prime}T_{n\ell}u_{n\ell}$, where $T_{nk}$ are either permutation, sign-change or orthogonal matrices, depending on the type of SA. We next show that, for every $k\neq\ell$, the quantity $f_{n,k\ell}:=d_{nk}d_{n\ell}g_{n,k\ell}/\sqrt{n}$ is stochastically bounded.
For $\mathcal{P}$-SA, this is seen by writing
$\sqrt{n}f_{n,k\ell}=$ $e_{k}^{\prime}D_{n}U_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}U_{n}D_{n}e_{\ell}=e_{k}^{\prime}V_{n}^{\prime}X_{n}^{\prime}H_{n}P_{nk}^{\prime}P_{n\ell}H_{n}X_{n}V_{n}e_{\ell}$
$=$ $e_{k}^{\prime}V_{n}^{\prime}(X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n})V_{n}e_{\ell}-ne_{k}^{\prime}V_{n}^{\prime}(\bar{x}_{n}\bar{x}_{n}^{\prime})V_{n}e_{\ell},$
which shows that
$f_{n,k\ell}=$ $\frac{1}{\sqrt{n}}e_{k}^{\prime}V_{n}^{\prime}(X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n})V_{n}e_{\ell}+o_{p}(1)=\frac{1}{\sqrt{n}}x_{n,k}^{\prime}P_{nk}^{\prime}P_{n\ell}x_{n,\ell}+o_{p}(1),$
where $x_{n,k}$ is the $k$th column of $X_{n}$. For the second equality, we have used the fact that $\frac{1}{\sqrt{n}}X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n}=\mathcal{O}_{p}(1)$ which holds immediately for the off-diagonal elements by the independence of $x_{n,k}$ and $x_{n,\ell}$. For the diagonal elements, the stochastic boundedness is equivalent to showing that $x_{n}^{\prime}P_{n}x_{n}/\sqrt{n}=\mathcal{O}_{p}(1)$ when $x_{n}\sim\mathcal{N}_{n}(0,I_{n})$ and $P_{n}$ is a uniformly random $n\times n$ permutation matrix. Since $\mathrm{E}(x_{n}^{\prime}P_{n}x_{n})=0$, by Theorem 14.4-1 in [2] we have
@@
Now, $\mathrm{E}(x_{n}x_{n}^{\prime}P_{n}x_{n}x_{n}^{\prime}|P_{n})=\mathrm{tr}(P_{n})I_{n}+P_{n}+P_{n}^{\prime}$, giving, using Lemma 1, that $\mathrm{Var}(x_{n}^{\prime}P_{n}x_{n})=\mathrm{E}[\{\mathrm{tr}(P_{n})\}^{2}]+\mathrm{E}\{\mathrm{tr}(P_{n}^{2})\}+n=n+4$. Plugging in to (6), the stochastic boundedness of $X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n}/\sqrt{n}$ and, consequently, that of $f_{n,k\ell}$ now follows.
For $\mathcal{J}$-SA, the reasoning is similar but we need the extra facts that $\frac{1}{n}X_{n}^{\prime}J_{nk}J_{n\ell}1_{n}=o_{p}(1)$ and $\frac{1}{n}1_{n}^{\prime}J_{nk}J_{n\ell}1_{n}=o_{p}(1)$ where the former holds because $J_{n\ell}J_{nk}X_{n}\sim X_{n}$. Furthermore, instead of Lemma 1, we use $\mathrm{tr}(J_{n}^{2})=n$ and $\mathrm{E}[\{\mathrm{tr}(J_{n})\}^{2}]=n$.
@@
What remains now is to show that $b_{n,kk},b_{n,k\ell}$, $k,\ell=1,\ldots,p$, $k\neq\ell$, have a limiting joint normal distribution with the desired covariance matrix. We begin by computing the covariance matrix.
@@
For $\mathcal{P}$-SA, direct computation gives $\mathrm{Var}(b_{n,kk})=2\lambda_{k}^{2}$ and $\mathrm{Var}(b_{n,k\ell})=2\lambda_{k}\lambda_{\ell}\{1+(1/n)\mathrm{E}\mathrm{tr}(P_{nk}^{\prime}P_{n\ell})\}=2(1+1/n)\lambda_{k}\lambda_{\ell}=2\lambda_{k}\lambda_{\ell}+o(1)$, where we have used the fact that $\mathrm{E}(P_{nk})=1_{n}1_{n}^{\prime}/n$. Similarly, we obtain the following non-zero covariances (rest of the covariances are zero): $\mathrm{Cov}(b_{n,k\ell},b_{n,\ell k})=\mathrm{Var}(b_{n,k\ell})=2\lambda_{k}\lambda_{\ell}+o(1)$. Analogous computation reveals that exactly the same asymptotic variances and covariances are obtained also for the two other forms of SA. Consequently, in each case, the limiting covariance matrix of $\sqrt{n}\mathrm{vec}(S_{n,\mathcal{P}}-\Lambda)$ is $(\Lambda^{1/2}\otimes\Lambda^{1/2})[2I_{p^{2}}+2K_{p,p}-2\mathrm{diag}\{\mathrm{vec}(I_{p})\}](\Lambda^{1/2}\otimes\Lambda^{1/2})$. The form for a general normal distribution now follows with equivariance arguments as in the proof of Theorem 2.1. That the moments of the limiting distribution actually are the limits of the moments we computed earlier, is guaranteed by showing that $b_{n,kk}$ and $b_{n,k\ell}$ are uniformly integrable. E.g., for $b_{n,k\ell}$, this can be done by denoting $A:=(1/\sqrt{n})x_{n,k}^{\prime}x_{n,\ell}$ and $B:=(1/\sqrt{n})x_{n,k}^{\prime}T_{nk}^{\prime}T_{n\ell}x_{n,\ell}$, using Young’s inequality to bound $\mathrm{E}\{(A+B)^{4}\}\leq a\mathrm{E}(A^{4})+b\mathrm{E}(B^{4})$ for some scalars $a,b\in\mathbb{R}$ and observing that $T_{nk}x_{n,k}\sim\mathcal{N}_{n}(0,I_{n})$ for each form of SA. The boundedness of $\mathrm{E}(A^{4})$ and $\mathrm{E}(B^{4})$ now follows from the moments of the Gaussian distribution.
Next, we still show that the limiting normal distribution actually exists. By the equivariance properties of the multivariate normal distribution, it is sufficient to consider only the case $\Lambda=I_{p}$ for this.
@@
We begin with $\mathcal{O}$-SA. Letting $Q_{n1},\ldots,Q_{np}$ be random orthogonal matrices uniform from the Haar distribution (and independent of all our other random variables), we have $x_{n,k}\sim Q_{nk}x_{n,k}$. Hence, (7) and (8) can be written as $b_{n,kk}=\sqrt{n}\{(1/n)x_{n,k}^{\prime}x_{n,k}-1\}+o_{p}(1)$ and $b_{n,k\ell}=(1/\sqrt{n})x_{n,k}^{\prime}(Q_{nk}^{\prime}Q_{n\ell}+R_{nk}^{\prime}R_{n\ell})x_{n,\ell}+o_{p}(1)$ where $R_{nk}:=O_{nk}Q_{nk}$ is Haar-uniformly distributed orthogonal matrix and independent of $Q_{nk}$. We then decompose $x_{n,k}=\chi_{nk}u_{nk}$ where $\chi_{nk}\sim\chi_{n}$ and $u_{nk}$ is uniform on the unit sphere and let $\chi_{nk1},\chi_{nk2}$ be independent $\chi_{n}$-random variables. Using these we write the elements $b_{n,kk}$ and $b_{n,k\ell}$ as
$b_{n,kk}$ $=\sqrt{n}\left(\frac{1}{n}(\chi_{nk}u_{nk})^{\prime}(\chi_{nk}u_{nk})-1\right)+o_{p}(1)$
@@
For $\mathcal{J}$-SA, the joint limiting normality follows directly from the standard multivariate CLT, once we observe that the transformations $J_{n1},\ldots,J_{np}$ do not “mix” the observation sequences and retain their i.i.d. nature.
@@
where we use Lemma 1, taking care of the third claim. Then, as $X_{1}$ equals $1/\sqrt{n}$ times a random variable with all moments finite, $X_{1}=(1/\sqrt{n})\{a_{11}(x_{1}^{2}-1)+a_{12}x_{1}(y_{1}+y_{\sigma(i)})+a_{22}(y_{1}^{2}-1)\}$, the first two claims in (9) also trivially hold. This conludes the proof in the case $p=2$ and the general case is handled exactly analogously.

### v2

# Katariina Perkonoja 1122 $0$-0002-3812-0871    Joni Virta 11 $0$-0002-2150-2769

@@ 1 Introduction
Figure 1: Illustration of the effects of $\mathcal{P}$-SA (blue), $\mathcal{J}$-SA (orange) and $\mathcal{O}$-SA (yellow) to the original data (green/darkest), with numerical labels corresponding to the row indices of original data, in a scenario with $n=10$ and $p=2$. $\mathcal{P}$-SA and $\mathcal{J}$-SA occasionally result in unwanted overlap, wherein a row in the anonymized data matches another row in the original dataset. In the case of $\mathcal{J}$-SA, this occurrence is due to coincidental alignment of signs with those in the original singular vectors, causing the observation to be duplicated (overlap of same indices), whereas for $\mathcal{P}$-SA, random permutation of values may cause another row to align with one in the original dataset (overlap of different indices). Conversely, $\mathcal{O}$-SA induces arbitrary rotations in the spectral space that prevent exact matches altogether with probability 1.
As our second main contribution, we investigate the earlier observation, stated also by Lasko and Vinterbo [9], that $\mathcal{P}$-SA approximately preserves the means, variances, covariances, and linear correlations of the original data $X_{n}$. In this work, we quantify this statement in an asymptotic framework. While SA itself makes no distributional assumptions and is applicable to any continuous data, we derive our results, for simplicity, under the assumption that the rows of $X_{n}$ are sampled i.i.d. from the normal distribution $\mathcal{N}_{p}(\mu,\Sigma)$ for some fixed $\mu\in\mathbb{R}^{p}$ and positive definite covariance matrix $\Sigma$. Extensions to other distributional assumptions are investigated through simulations in Section 3 and discussed further in Section 4. As the multivariate normal distribution is entirely determined by its means and covariances, evaluating the resemblance of anonymized data with respect to these moments is equivalent to assessing general data utility in the present scenario.
@@ 2 Asymptotic utility
Theorem 2.2 offers us two insights: (a) All three forms of SA are equally efficient in covariance estimation. (b) Assume next that $O=I_{p}$ in the eigendecomposition $\Sigma=O\Lambda O^{\prime}$ (since the rotation $O$ only fixes the coordinate system in which the data is observed, this choice does not limit our interpretations). Then, for instance when $p=2$, the limiting covariance matrices of $S_{n}$ and $S_{n,\mathcal{A}}$ take the forms:
@@ 3 Simulations
Figure 2: The relative error of the empirical covariance matrices of sample mean (solid line) and sample covariance (dotted line) compared to their asymptotic covariance matrices for $\mathcal{P}$-SA (blue), $\mathcal{J}$-SA (orange), and $\mathcal{O}$-SA (yellow), when original data (green/darkest) is sampled from a normal distribution meeting the Assumption 1. Sample size is presented on a logarithmic scale.
@@ 3.1 Finite-sample utility
Figure 3: The relative error of the empirical covariance matrices of sample mean (solid line) and sample covariance (dotted line) compared to their asymptotic covariance matrices for $\mathcal{P}$-SA (blue), $\mathcal{J}$-SA (orange), and $\mathcal{O}$-SA (yellow), when original data (green/darkest) is sampled from a normal distribution violating the Assumption 1. Sample size is presented on a logarithmic scale.
@@ 4 Conclusion
Based on our findings with finite data, none of the three methods demonstrated clear superiority in utility convergence. However, concerning privacy, $\mathcal{O}$-SA outperformed $\mathcal{P}$-SA and $\mathcal{J}$-SA by never generating identical records. Hence, we advocate for employing $\mathcal{O}$-SA for data anonymization, notwithstanding its computational intensity compared to $\mathcal{P}$-SA and $\mathcal{J}$-SA. For a dataset size of $n=1000$ and $p=6$, the runtime for $\mathcal{O}$-SA is on average 2.44 seconds, while for $\mathcal{P}$-SA and $\mathcal{J}$-SA, it is 0.002 and 0.001 seconds, respectively (using AMD Ryzen 5 5600X). The complexity of SVD is the same as for principal component analysis, meaning that the method scales well. In practical scenarios, data anonymization is typically a one-time operation, mitigating concerns over computational overhead, and all SA algorithms can be applied to larger datasets (both in terms of $n$ and $p$) provided the data is continuous. However, ensuring asymptotic utility (Section 2) necessitates additional assumptions. Should computational efficiency become paramount, $\mathcal{P}$-SA emerges as the secondary choice (no matches in our simulations with $p=6$). In case privacy is given secondary priority, then $\mathcal{P}$-SA offers superior mean estimation efficiency to its two competitors (Theorem 2.1) and should be used. Nonetheless, further research is warranted to evaluate these SA algorithms using empirical real-world data and utility scenarios.
Our simulations showed that Assumption 1 cannot be dispensed with. SA is based on perturbing the singular vectors of $H_{n}X_{n}$ and, intuitively, the role of Assumption 1 is to ensure that these vectors are asymptotically identifiable. Similarly, application to Poisson-distributed data revealed that our conclusions do not directly extend outside of normality. However, we expect that equivalent results could be obtained also under other distributional assumptions, with suitable techniques of proof. And while we chose to derive our theoretical results under the normality assumption due to the ubiquitousness of the Gaussian distribution, the SA-method itself remains applicable also outside of normal data.
A prospective research direction would be to conduct analogous studies for competing methods, enabling utility and privacy comparisons between the methods. E.g., data shuffling [11] and correlated noise addition [15] have been shown to preserve second-order statistics. Unlike SA, data shuffling and noise addition use distributional assumptions as part of their data generation and their impact on the methods’ utility could be quantified with results such as Theorem 2.1 and 2.2. For categorical data, Post-randomization Method (PRAM) leads to unbiased estimates of univariate moments, see, e.g. [15], and also seems applicable to asymptotic analysis. However, PRAM operates on a single variable at a time and more elaborate techniques are required to preserve the dependency structures between multiple categorical variables. Finally, investigating modification constraints, such as selectively altering the first $j$ columns of $U_{n}$, presents a compelling research area, especially for high-dimensional datasets.
@@ 4.0.1 Acknowledgements
The study was supported by the Finnish Cultural Foundation (grant 00220801) and the Research Council of Finland (grants 347501 and 353769). The authors would like to thank two anonymous reviewers for their valuable comments.
@@ Appendix 0.A Proofs of technical results
For $\mathcal{O}$-SA, the proof is similar to $\mathcal{J}$-SA, and we point out only the differences next. To see that $X_{n}^{\prime}H_{n}O_{nk}1_{n}/\sqrt{n}$ is $\mathcal{O}_{p}(1)$, we observe that $O_{nk}1_{n}\sim\sqrt{n}O_{nk}e_{1}\sim\sqrt{n}u_{n}$, where $u_{n}$ is uniform on the unit sphere $\mathbb{S}^{n-1}$. Moreover, letting $\chi_{n}$ denote a random variable obeying a $\chi$-distribution with $n$ degrees of freedom that is independent of $u_{n}$, we have $b_{n}:=\chi_{n}u_{n}\sim\mathcal{N}_{n}(0,I_{n})$ and $\chi_{n}/\sqrt{n}\rightarrow_{p}1$. Consequently, denoting a generic column of $X_{n}$ by $x_{n}\sim\mathcal{N}_{n}(0,\lambda I_{n})$, we have $(1/\sqrt{n})x_{n}^{\prime}H_{n}O_{nk}1_{n}\sim x_{n}^{\prime}H_{n}u_{n}=x_{n}^{\prime}u_{n}-\bar{x}_{n}1_{n}^{\prime}u_{n}=b_{n}^{-1}(1/\sqrt{n})x_{n}^{\prime}z_{n}-\sqrt{n}\bar{x}_{n}\bar{z}_{n}b_{n}^{-1}$, showing that $X_{n}^{\prime}H_{n}O_{nk}1_{n}/\sqrt{n}$ is stochastically bounded. Hence,
$(\sqrt{n}\bar{y}_{n})_{k}=(\sqrt{n}\bar{x}_{n})_{k}+b_{n}^{-1}\frac{1}{\sqrt{n}}e_{k}^{\prime}X_{n}^{\prime}z_{n}+o_{p}(1)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}x_{n,ik}(1+z_{n,i})+o_{p}(1).$
@@
as $\sqrt{n}\bar{y}_{n}=\mathcal{O}_{p}(1)$. From Theorem 2.1 we have $V_{n}D_{n}U_{n0}^{\prime}1_{n}\bar{x}_{n}/\sqrt{n}=o_{p}(1)$, giving
@@
where $G_{n}\in\mathbb{R}^{p\times p}$ has $g_{n,kk}=0$, $g_{n,k\ell}=u_{nk}^{\prime}T_{nk}^{\prime}T_{n\ell}u_{n\ell}$, where $T_{nk}$ are either permutation, sign-change or orthogonal matrices, depending on the type of SA. We next show that, for $k\neq\ell$, the quantity $f_{n,k\ell}:=d_{nk}d_{n\ell}g_{n,k\ell}/\sqrt{n}$ is $\mathcal{O}_{p}(1)$.
For $\mathcal{P}$-SA, this is seen by writing $\sqrt{n}f_{n,k\ell}=e_{k}^{\prime}D_{n}U_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}U_{n}D_{n}e_{\ell}=e_{k}^{\prime}V_{n}^{\prime}X_{n}^{\prime}H_{n}P_{nk}^{\prime}P_{n\ell}H_{n}X_{n}V_{n}e_{\ell}=e_{k}^{\prime}V_{n}^{\prime}(X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n})V_{n}e_{\ell}-ne_{k}^{\prime}V_{n}^{\prime}(\bar{x}_{n}\bar{x}_{n}^{\prime})V_{n}e_{\ell}$, giving $f_{n,k\ell}=n^{-1/2}e_{k}^{\prime}V_{n}^{\prime}(X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n})V_{n}e_{\ell}+o_{p}(1)=n^{-1/2}x_{n,k}^{\prime}P_{nk}^{\prime}P_{n\ell}x_{n,\ell}+o_{p}(1)$, where $x_{n,k}$ is the $k$th column of $X_{n}$. For the second equality, we have used $\frac{1}{\sqrt{n}}X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n}=\mathcal{O}_{p}(1)$ which holds for the off-diagonal elements by the independence of $x_{n,k}$ and $x_{n,\ell}$. For the diagonal elements, the boundedness is equivalent to $x_{n}^{\prime}P_{n}x_{n}/\sqrt{n}=\mathcal{O}_{p}(1)$ when $x_{n}\sim\mathcal{N}_{n}(0,I_{n})$ and $P_{n}$ is a uniformly random $n\times n$ permutation matrix. Since $\mathrm{E}(x_{n}^{\prime}P_{n}x_{n})=0$, by [2, Theorem 14.4-1],
@@
Now, $\mathrm{E}(x_{n}x_{n}^{\prime}P_{n}x_{n}x_{n}^{\prime}|P_{n})=\mathrm{tr}(P_{n})I_{n}+P_{n}+P_{n}^{\prime}$, giving, using Lemma 1, that $\mathrm{Var}(x_{n}^{\prime}P_{n}x_{n})=\mathrm{E}[\{\mathrm{tr}(P_{n})\}^{2}]+\mathrm{E}\{\mathrm{tr}(P_{n}^{2})\}+n=n+4$. Plugging in to (6), the stochastic boundedness of $X_{n}^{\prime}P_{nk}^{\prime}P_{n\ell}X_{n}/\sqrt{n}$ and $f_{n,k\ell}$ now follows.
For $\mathcal{J}$-SA, the reasoning is similar but uses the facts that $\frac{1}{n}X_{n}^{\prime}J_{nk}J_{n\ell}1_{n}=o_{p}(1)$ and $\frac{1}{n}1_{n}^{\prime}J_{nk}J_{n\ell}1_{n}=o_{p}(1)$ where the former holds because $J_{n\ell}J_{nk}X_{n}\sim X_{n}$. Furthermore, instead of Lemma 1, we use $\mathrm{tr}(J_{n}^{2})=n$ and $\mathrm{E}[\{\mathrm{tr}(J_{n})\}^{2}]=n$.
@@
What remains now is to show that $b_{n,kk},b_{n,k\ell}$, $k,\ell=1,\ldots,p$, $k\neq\ell$, have a limiting joint normal distribution with the desired covariance matrix. We begin by computing the covariance matrix. For $\mathcal{P}$-SA, direct computation gives $\mathrm{Var}(b_{n,kk})=2\lambda_{k}^{2}$ and $\mathrm{Var}(b_{n,k\ell})=2\lambda_{k}\lambda_{\ell}\{1+(1/n)\mathrm{E}\mathrm{tr}(P_{nk}^{\prime}P_{n\ell})\}=2(1+1/n)\lambda_{k}\lambda_{\ell}=2\lambda_{k}\lambda_{\ell}+o(1)$, where we have used the fact that $\mathrm{E}(P_{nk})=1_{n}1_{n}^{\prime}/n$. Similarly, we obtain the following non-zero covariances (rest of the covariances are zero): $\mathrm{Cov}(b_{n,k\ell},b_{n,\ell k})=\mathrm{Var}(b_{n,k\ell})=2\lambda_{k}\lambda_{\ell}+o(1)$. Analogous computation reveals that the same asymptotic variances and covariances are obtained also for $\mathcal{J}$-SA and $\mathcal{O}$-SA. Consequently, in each case, the limiting covariance matrix of $\sqrt{n}\mathrm{vec}(S_{n,\mathcal{P}}-\Lambda)$ is $(\Lambda^{1/2}\otimes\Lambda^{1/2})[2I_{p^{2}}+2K_{p,p}-2\mathrm{diag}\{\mathrm{vec}(I_{p})\}](\Lambda^{1/2}\otimes\Lambda^{1/2})$. The form for a general normal distribution now follows with equivariance arguments as in the proof of Theorem 2.1. That the moments of the limiting distribution actually are the limits of the moments we computed earlier, is guaranteed by showing that $b_{n,kk}$ and $b_{n,k\ell}$ are uniformly integrable. E.g., for $b_{n,k\ell}$, this can be done by denoting $A:=(1/\sqrt{n})x_{n,k}^{\prime}x_{n,\ell}$ and $B:=(1/\sqrt{n})x_{n,k}^{\prime}T_{nk}^{\prime}T_{n\ell}x_{n,\ell}$, using Young’s inequality to bound $\mathrm{E}\{(A+B)^{4}\}\leq a\mathrm{E}(A^{4})+b\mathrm{E}(B^{4})$ for some scalars $a,b\in\mathbb{R}$ and observing that $T_{nk}x_{n,k}\sim\mathcal{N}_{n}(0,I_{n})$ for each form of SA. The boundedness of $\mathrm{E}(A^{4})$ and $\mathrm{E}(B^{4})$ now follows from the moments of the Gaussian distribution.
@@
Next, we show that the limiting distribution exists. By the equivariance properties of the multivariate normal, it is sufficient to consider only the case $\Lambda=I_{p}$. We begin with $\mathcal{O}$-SA. Letting $Q_{n1},\ldots,Q_{np}$ be random orthogonal matrices uniform from the Haar distribution (and independent of all our other random variables), we have $x_{n,k}\sim Q_{nk}x_{n,k}$. Hence, (7) and (8) can be written as $b_{n,kk}=\sqrt{n}\{(1/n)x_{n,k}^{\prime}x_{n,k}-1\}+o_{p}(1)$ and $b_{n,k\ell}=(1/\sqrt{n})x_{n,k}^{\prime}(Q_{nk}^{\prime}Q_{n\ell}+R_{nk}^{\prime}R_{n\ell})x_{n,\ell}+o_{p}(1)$ where $R_{nk}:=O_{nk}Q_{nk}$ is Haar-uniformly distributed orthogonal matrix and independent of $Q_{nk}$. We then decompose $x_{n,k}=\chi_{nk}u_{nk}$ where $\chi_{nk}\sim\chi_{n}$ and $u_{nk}$ is uniform on the unit sphere and let $\chi_{nk1},\chi_{nk2}$ be independent $\chi_{n}$-random variables. Using these we write $b_{n,kk}$ and $b_{n,k\ell}$ as $b_{n,kk}=\sqrt{n}\{(1/n)(\chi_{nk}u_{nk})^{\prime}(\chi_{nk}u_{nk})-1\}+o_{p}(1)$ and as
@@
For $\mathcal{J}$-SA, joint limiting normality follows from the multivariate CLT, as $J_{n1},\ldots,J_{np}$ do not “mix” the observations and retain their i.i.d. nature.
@@
where we use Lemma 1, taking care of the third claim. Then, as $X_{1}$ equals $1/\sqrt{n}$ times a random variable with all moments finite, $X_{1}=(1/\sqrt{n})\{a_{11}(x_{1}^{2}-1)+a_{12}x_{1}(y_{1}+y_{\sigma(i)})+a_{22}(y_{1}^{2}-1)\}$, the first two claims in (9) also trivially hold. This conludes the proof in the case $p=2$ and the general case is handled analogously.